Thursday, April 14, 2011

Day 16, 04/14/2011 -- G.I. Joe kung fu gripper

Today, we again attacked the problem of the HWP not rotating. We tried a few different things, like hitting the suspect gripper with tons of current and tilting the cryostat to different angles. Still, despite our best efforts, we have yet to find a solution. We do feel we have a better idea of which gripper is not working, and there are some things we can try tomorrow when we go in.

Franky has finished setting up the BROs; tomorrow we plan on connecting cables and cycling the fridges so we can start doing detector work.


Wednesday, April 13, 2011

Day 15, 04/13/2011 -- Something is rotten in the state of Texas

Rotten is perhaps too strong a word. Basically, when you do experimental physics, things often don't work. It's part of the simultaneous charm and frustration of doing new things.

In this case, we are having trouble with the "grippers", a set of three motor-controlled jaws that hold the half-wave-plate in place as the cryostat cools down. Despite repeated testing in another cryostat and Jeff's careful assembly, we have yet to succeed in opening the grippers and releasing the HWP. We still have a few things to try tomorrow, so hope is far from lost. Even if the HWP isn't free to rotate, we still have plenty of other tests that need to be done, both without and eventually with the gondola, so we are forging ahead.

While the UMN crew tackled the gripper problem, Franky kept himself busy with continuing to assemble and populate the BROs. At this point, we're just waiting for the detectors to be cold enough that we can start connecting cables and doing the baseline checkout tests.

After Jeff and I went in at 4:30 AM for a helium transfer, the boiloff rate is now slowing down and will likely last through the night -- a good sign that we are close to being able to run the fridges and detectors.


Tuesday, April 12, 2011

Day 14, 04/12/2011 -- At last, helium.

With the arrival of Franky from McGill yesterday, today he set about setting up the bolometer readout crates (hereafter: BROs). After first assembling the final crate, he identified and fixed a power issue related to a bent pin that caused us to damage 2 (very expensive) readout boards back in Minnesota. The main problem he and Kyle were having was failure to communicate -- no, not because of his Quebecois accent, but because they were having network problems between the computers and the BROs. In the end it turned out to be something misconfigured in one of our network switches, but it took a half day to figure out what the problem was and how to fix it.
In the meantime, Kyle, Jeff, and I started thinking about some of the upcoming calibration experiments -- how to mount a source on the water tower and be able to aim it, how to suspend a styrofoam bucket with LN2 above the cryostat window, etc. One of the outcomes of this discussion was a nice little mount based on a benchtop optical mount, a spare hexapod clevis, and a piece of scrap aluminum. This mount, an inverted altitude-azimuth mount destined for holding our millimeter-wave source atop the water tower, will allow us to map the beams of the telescope once the gondola gets here.

But the main advance of the day was helium, and lots of it. We finally got our gas regulator from Minnesota and could proceed with cooling down the cryostat with liquid helium. After pumping out most of the LN2 from the helium tank and then boiling off the little that remained with warm nitrogen gas, we then proceeded to transfer liquid helium out of one of the large 500L storage dewars we ordered. In the end we used about 250L of helium to cool and completely fill the EBEX cryostat's 130L helium tank. Because the optics box and focal planes inside are isolated from the 4K cryostat shell, they still need to cool down, and as they do, they boil off massive amounts of liquid helium. Thus, for the next few days, we'll be doing multiple fills a day, at all hours of the day or night -- Jeff and I are scheduled for a fill at about 3AM, and there will probably be another stupidly early/late fill tomorrow as well. Ah, the joys of having a big cryostat!

Pictures, again, finally:

Sunday, April 10, 2011

Day 12, 04/10/11 -- Tippecanoe and Tyler Too

Today was a slow day. With the cryostat filled with liquid nitrogen and the internals slowly cooling, there's not much that requires immediate attention. Since it's Sunday we went into the high bay only for a short period of time today -- long enough to check the liquid nitrogen levels (and refill as needed) and for Jeff to open and close the internal mechanical heatswitch to try and optimize its thermal contact.

We then drove to Tyler, TX to see what it was all about. A city of over 80,000 people, we figured it must have something worthwhile, right? Turns out we were right: Jeff found that they have a zoo and a historical aircraft and aviation museum. We went to the Caldwell Zoo first, and were treated with a surprisingly nicely maintained and well-curated zoo, with all manner of big cats (lions, white tigers, leopards, and cheetahs), a couple of black bears, and some large mammals from both the US and Africa (rhinos, giraffes, and elephants) along with the requisite reptile area (which focused mostly on species native to the area) and a nice contingent of water and tropical birds.

The aviation museum was also quite a find. Located at the Tyler airport, they have a nice collection of aviation-related memorabilia from the entire history of the US armed forces as well as a nice collection of real (but mostly engineless) aircraft sitting on the tarmac (an F-101, F-105, F-111, and an F-4 Phantom come to mind immediately out of the 10 or so aircraft they had). They also had knowledgeable and helpful docents on hand to answer questions about the aircraft, which was a nice bonus. And all for $5! Highly recommended if you're ever in Tyler.

The only pictures from today are some cell phone pictures of pelicans, so I'm not going to bother. Tomorrow we're back in the high bay, and are planning our first of many liquid helium transfers!

Saturday, April 9, 2011

Day 11, 04/09/11 -- The Birth of Cool

We spent a half day in the high bay today getting the cryostat ready to cool with liquid nitrogen. Jeff and I worked on getting electronics boards mounted to the bottom of the cryostat as well as connecting their cables inside the metallic 'can' that sits at the bottom of the cryostat. Kyle meanwhile worked on getting the cryostat housekeeping (temperature sensors, heaters, and current monitors) up and running.

After installing cables and the RF can, we flipped the cryostat upright to begin the cooldown. The liquid nitrogen tank obviously holds liquid nitrogen, but we pre-cool the liquid helium tank with liquid nitrogen as well. Liquid nitrogen has a much higher heat capacity and is much less expensive than liquid helium, so we use LN2 to get us from 300K to 77K first, and then fill with LHe. That process went quite smoothly despite (or perhaps because?) the cryogen vendor gave us high pressure (250 psi!) LN2 tanks instead of the 22psi ones we requested. At any rate, it worked, and Kyle got the housekeeping system up and running after some minor wiring tweaks to account for some changes inside the cryostat.

After leaving the high bay, Kyle headed back to our lodgings while Jeff and I decided to take a bike ride. It turns out that not only is the Palestine area quite pretty as we've noticed since being here, with lots of trees and green fields, but the Texas drivers were surprisingly accommodating of two guys riding skinny-tired bikes on the farm roads and state highways. On the smaller roads, drivers without fail gave us a wide berth, and we were not once honked at from someone behind us (one guy in town coming the opposite way honked at us for some unknown reason...or perhaps he was honking at someone else). Also, Palestine's hills are surprisingly painful for someone (me) who hasn't done any real riding since last October in Minneapolis (which is largely flat). Still, the 14 miles we did today was a good start -- hopefully we'll have a chance to do some longer rides in the future.

The few pictures today were taken with my cell phone, as my DSLR's memory card was left in the apartment. Sorry.

Day 10, 04/08/11 -- Leak free

Whoops, sorry for the late update.

Because we had a late night on Thursday and have been working pretty much all day every day, we took our first (mostly) day off. The only thing we did in the high bay was leak-checking the cryostat. This was somewhat complicated by the fact that we forgot to bring an inert gas regulator with us to Palestine. We managed to borrow one from the other science group (CREAM) that is in the high bay adjacent to ours.

Planning to attach cables, flip the cryostat, and fill with LN2 today.

No photos from Friday because there wasn't anything interesting to take pictures of...

Thursday, April 7, 2011

Day 9, 04/07/11 -- Closing Time

After a long day, we got the cryostat closed and pumping! It's late and I want to go to bed so no detailed update tonight, but everything went pretty smoothly. Hopefully it won't have a leak -- fingers crossed!


Wednesday, April 6, 2011

Day 8 , 04/06/11 -- Just talkin' 'bout shaft

Another day, another few steps closer to getting the cryostat closed and pumping. We pulled the instrument out of the cryostat for the final time so that Jeff could finish installing the parts for the half-wave-plate assembly.

Jeff spent a large fraction of the day inside the cryostat. We had a little scare early in the day as Jeff was installing part of the HWP drive shaft assembly -- it got dropped on the floor, bending two of the flexible bellows in the process. After a brief moment of panic, Jeff realized he actually had spares for the parts in his parts bin, so after a quick replacement we were back on our way. After that, the wiring for the HWP system was installed and tested. The last piece of the puzzle was measuring positions of the angle encoder components on the HWP assembly so we can later infer the HWP's angle relative to the rest of the cryostat and, by extension, the sky.

I spent much of the day working on getting our optical filter, made of porous Teflon, ready to mount. This filter, which helps protect the 1 Kelvin optics from the thermal radiation from the relatively warm (~30K) rotating HWP, is a new addition for this cryostat run.

Kyle spent a lot of the day working on getting our computer network set up, which he did with the help of Milligan back in Minnesota and the CSBF networking person. He also started getting our vacuum leak checker, which arrived today from Minnesota, set up. Kyle and I both helped Jeff where we could, as he had the most difficult work for today.


Tuesday, April 5, 2011

Day 7, 04/05/11 -- Tape'n'tape

Today was a productive day in the high bay. Jeff started installing the half-wave-plate motor driveshaft assembly and made a lot of progress in putting together the 'table' assembly that holds the entire HWP system.

Kyle and I measured the polarizing grid orientation relative to the rest of the optics and then we started installing the shielding that protects the detectors from seeing relatively warm 4.2K radiation from the cryostat walls.

This process involves a lot of tape. Aluminum foil tape, to be precise. Sure, the shielding is screwed down, but all of the little gaps between parts need to be closed off with tape. It's a pretty tedious job, as there are a lot of seams, and some of the seams are tough to reach or have non-trivial geometry that makes them difficult to cover with tape. Still, we do our best, and the process takes a few hours to do well.

Once the instrument was all taped up, we got the CSBF guys to come in and crane the instrument over and into the cryostat. Once we got it in place, Kyle spent the rest of the day aligning the last optical element in the cryostat, the so-called "field" lens. In the meantime, I helped Jeff as much as I could, doing little things that needed to be done while he fussed with tiny bearings for the driveshaft assembly.

Tomorrow will be devoted mostly to HWP system installation (finishing the driveshaft, installing/connecting wires, and then installing the HWP system itself). The hope is we can close up and start pumping on the cryostat on Thursday.


Monday, April 4, 2011

Day 6, 04/05/11 -- Lens Alignment: Chaotic Neutral?

Today was mostly devoted to finishing the alignment process Kyle started in the optics box. We fixed our mechanical overconstraint as mentioned yesterday, but we were still getting unreliable results from our lens alignment procedure. We finally realized that we actually had to rethink part of the process: we were trying to align to something that had too many interdependent parameters. Once we picked a different reference point to align to, the process went more smoothly. Kyle also installed and aligned the polarizing grid in the optics box, so now all of the 1 Kelvin optics are done.

In the meantime, Jeff continued to work on the cold wiring and assembly of the half-wave-plate assembly. One issue that came up was trying to figure out how best to understand the position of the HWP relative to the rest of the cryostat. The HWP rotates, but at any given point in time you need to know its orientation to some known coordinate system so you can figure out the polarization on the sky. It turns out that this isn't a trivial task, but we came up with a procedure for measuring the position of the HWP assembly relative to the outside of the cryostat that should work. We spent a lot of time verifying that this procedure was robust by making many test measurements and checking their repeatability.

Last, I appear to have hurt myself. I must have done something to my ankle while running the past couple of days (Friday and Sunday) and I'm now walking around (slowly) with a noticeable limp. I'm sure it wasn't helped by having to climb up and down ladders in order to make measurements. Hopefully with ice and compression, it will get better by itself in a couple of days...fingers crossed!


Sunday, April 3, 2011

Day 5, 04/05/11 -- Lazy Sunday

Ok, it wasn't really a "lazy sunday" -- we still went in to the high bay, but only after a proper breakfast and we "only" stayed 9 hours.

Kyle spent all day doing cold optics alignment. Normally, this wouldn't take all day, but for some reason Kyle was having a lot of trouble with one lens in the optics box. after scratching our heads for a bit, we realized that the alignment scheme was mechanically overconstrained, which we theorize was leading to strange twists in the lens mount structure. We have removed the overconstraint and are going to tackle the alignment again in the morning.

Meanwhile, I finished up replacing the connectors on Jeff's feedthrough wiring and potted them in epoxy (this time after sealing up the back of the connector with a silicone conformal coating to prevent a repeat of last time). Jeff meanwhile replaced components in the top optics stack of the cryostat. We have one new thermal filter from our collaborators at Cardiff and we hope to have a second in time for the closing of the cryostat. These filters are of a new design that should lead to lower helium boiloff in the cryostat -- always a good thing. Jeff also started work on replacing some half-wave-plate system wiring that lives at 4 Kelvin to account for changes made since the last time the HWP was inside the cryostat.

On a non-science note, Jeff and I have found that the big circular launch pad at CSBF makes for a nice running track. We're clearly not the first to think of this, since the road around the pad is already marked with distance indicators. But since I've decided that I'm going to try and improve my running while here in Palestine (in lieu of my usual physical activities), I decided to get a baseline today. I ran the 1.2 mile loop in 8:20, which corresponds to a 6:57 mile -- way faster than I thought I could run a mile! I'll keep you updated on how I improve (or not) throughout our time in Palestine.

No pictures today, since there was really nothing new worth taking pictures of compared to yesterday. Make sure to check out yesterday's pictures, though, if you missed them (since I added them this morning).

Saturday, April 2, 2011

Day 4, 04/02/11 -- Epoxy, for example

In case you hadn't realized, yesterday's post was a (lame) April Fool's joke. The Helium-10 fridge is fine, and everything is moving along.

Yesterday's update was basically more progress on all fronts. Kyle continued to put the instrument together, I finished the housekeeping wiring feedthrough wire replacement and made a new charcoal getter for the cryostat while Jeff finished soldering his wiring. Like much of our wiring, Jeff potted the connectors, which means he encapsulated the wires at the back of the connectors with epoxy (in this case, Stycast 2850FT, an alumina-filled epoxy designed for relatively high thermal conductivity and a metal-like thermal expansion coefficient) in order to prevent wires from moving and shorting together. We typically use the 2850FT resin with a low-viscosity catalyst, which is great for getting the epoxy to flow between wires, around contacts, and into the nooks and crannies (sort of like an english muffin) of the mold. Unfortunately for Jeff, the connectors he used on one end of his wiring, for some reason, had somewhat loose contacts, and the epoxy flowed around them and into the front face of the connector, rendering it unusable.

Today we decided to cut off the bad connectors and replace them. Luckily, unlike most of the connectors in the experiment, these particular connectors (25-pin D-subminiature connectors, the ubiquitous DB-25) are common and we were able to get replacements from the local Radio Shack. These connectors seem to have much more firmly-embedded contacts, so we don't expect the same problem to occur.

On other fronts, Kyle continued with the instrument -- he finished up making all of the sub-Kelvin thermal connections and we started on the process of optics alignment. This is a tricky task with tight tolerances: 0.005" in translation and 0.1 degree in angle. In order to measure these distances, we use a neat tool called a Microscribe MX, which is an articulated arm used to measure locations of...basically anything. That process continues.

In the meantime, I worked on some internal cryostat wiring while Jeff removed the components in the so-called "optics stack" of the instrument, which holds various optical filters and components, in preparation for upcoming alignment tasks.

Unfortunately, I forgot my camera's memory card and card reader in the high bay, so I have no pictures from today (I only took a few anyway). You'll have to make do with the pictures from yesterday that I didn't upload. I'll edit tomorrow with the link to today's pictures.

Yesterday's pictures:

EDIT: Today's pictures:

Friday, April 1, 2011

Day 3, 04/01/11 -- Tragedy

This is not good.

We needed to move something out of the way (the 150 lb flat bottom lid used when transporting our cryostat) so we foolishly decided to use the crane ourselves. After we had lifted the lid and as we started moving it to the other end of the high bay, the tag line for controlling the load broke and the lid swung over and hit the instrument.

It ended up putting a crack in one of the thin-walled tubes in our so-called "helium-10" fridge, which subsequently lost all of its helium charge. On the plus side, it was only helium 4, which isn't nearly as expensive as helium 3. On the minus side, without helium 4, we can't cycle the fridge to get the detectors cold.

This is not a good way to start the month of April.